OpenStudy (anonymous):

(2x^2+10x+12)+(x+3) using synthetic division

OpenStudy (anonymous):

need help plzz

OpenStudy (anonymous):

ill give free medals

OpenStudy (dumbcow):

to use synthetic division, you must be dividing by (x-c) so in this case c = -3 -3 | 2 10 12 ______-6___-12___ 2 4 0 remainder is zero quotient is 2x+4

OpenStudy (anonymous):

so the answer is 2x+4 r0

OpenStudy (dumbcow):

you could have just factored here 2x^2+10+12 = 2(x+3)(x+2)

OpenStudy (anonymous):

can you help me with more plzzz

OpenStudy (dumbcow):

yes that is answer

OpenStudy (anonymous):

oh thank you so much

OpenStudy (anonymous):

i have a another one

OpenStudy (anonymous):

\[x ^{3}+4x ^{2}-3x-12 divided by x ^{2}-3\]

OpenStudy (dumbcow):

factor

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

you can use factor by grouping...im waiting for others to respond before i explain further

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lil confused

OpenStudy (anonymous):

\[i read the statement as divide x ^{3}+4x ^{2}-3x-12 by x ^{2} -3\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the way you set it up was confusing

OpenStudy (anonymous):

i used long division

OpenStudy (anonymous):

they told me to use synthetic division

OpenStudy (dumbcow):

thats polynomial division...have you learned that yet? again its not necessary here tho \[\frac{(x^{3}+4x^{2})+(-3x-12)}{x^{2}-3} = \frac{x^{2}(x+4)-3(x+4)}{x^{2}-3} = \frac{(x+4)(x^{2}-3)}{x^{2}-3} = x+4\]

OpenStudy (anonymous):

no not really

OpenStudy (dumbcow):

synthetic division wont work unless its in form x-c

OpenStudy (dumbcow):

@surjithayer , you made a mistake in your long division

OpenStudy (anonymous):

oh ok can you explain it to me then

OpenStudy (dumbcow):

i already did...i used factor by grouping to factor numerator ^^

OpenStudy (anonymous):

ohh lol ok

OpenStudy (anonymous):

\[x ^{2}-3=0,x ^{2}=3,x=\sqrt{3},-\sqrt{3}\]  Quoient =x+4, Remainder 0